ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 93, 1996

Acosta, M. D., F. J. Aguirre and R. Payá There is no bilinear Bishop–Phelps theorem

221

Aguirre, F. J. See M. D. Acosta

Beaulieu, R. A., A. Jensen and S. Jøndrup Towards a more general notion of Gelfand–Kirillov dimension

73

Benyash-Krivetz, V. V. See A. S. Rapinchuk

Berndt, J., F. Prüfer and L. Vanhecke Geodesic spheres and two-point homogeneous spaces

373

Chernousov, V. I. See A. S. Rapinchuk

Chothi, V., G. Everest and T. Ward Oriented local entropies for expansive actions by commuting automorphisms

281

Choucroun, F. M. Sous-groupes discrets des groupes 𝔭-adiques de rang un et arbres de Bruhat-Tits

195

Coelho, Z. and E. de Faria Limit laws of entrance times for homeomorphisms of the circle

93

de Faria, E. See Z. Coelho

Everest, G. See V. Chothi

Feit, W. Schur indices of characters of groups related to finite sporadic simple groups

229

Freud, R. and P. P. Pálfy On the possible number of elements of given order in a finite group

345

Gitik, M. and J. Witzany Consistency strength of the axiom of full reflection at large cardinals

113

Haran, D. and H. Völklein Galois groups over complete valued fields

9

Iranzo, M. J., G. Navarro and F. Pérez Monasor A conjecture on the number of conjugacy classes in a p-solvable group

185

Jensen, A. See R. A. Beaulieu

Jøndrup, S. See R. A. Beaulieu

Junco, A. del and D. J. Rudolph Residual behavior of induced maps 387 Katok, A. and J. Lewis Global rigidity results for lattice actions on

tori and new examples of volume-preserving actions

253

Kra, B. The conjugating map for commutative groups of circle diffeomorphisms

303

Lewis, J. See A. Katok

Liao, G.-F. and L.-Y. Wang Almost periodicity, chain recurrence and chaos

145

Lindenstrauss, A. Deformation retracts and the Hochschild homology of polynomial rings

317

Matous̆ek, J. On the distortion required for embedding finite metric spaces into normed spaces

333

Mekler, A. H., S. Shelah and O. Spinas The essentially free spectrum of a variety

1

Monterde, J. and O. A. Sánchez-Valenzuela The exterior derivative as a Killing vector field

157

Navarro, G. See M. J. Iranzo

Pálfy, P. P. See R. Freud

Pawlikowski, J. A characterization of strong measure zero sets  171

Payá, R. See M. D. Acosta

Pérez Monasor, F. See M. J. Iranzo

Prüfer, F. See J. Berndt

Quas, A. N. A C^{1} expanding map of the circle which is not weakmixing

359

Rapinchuk, A. S., V. V. Benyash-Krivetz and V. I. Chernousov Representation varieties of the fundamental groups of compact orientable surfaces

29

Rudolph, D. J. See A. del Junco

Sánchez-Valenzuela, O. A. See J. Monterde

Schneps, L. On Galois groups and their maximal 2-subgroups  125
Schramm, O. Conformal uniformization and packings  399

Shelah, S. See A. H. Mekler

Spinas, O. See A. H. Mekler

Vanhecke, L. See J. Berndt

Völklein, H. See D. Haran

Wang, L.-Y. See G.-F. Liao

Ward, T. See V. Chothi

Witzany, J. See M. Gitik

Xu, K. Nilpotent subgroups of self equivalences of torsion spaces  189